OpenStudy (anonymous):

1/3a = 3/10 (Solve the Equation) (Multiplying and Dividing Rational Numbers) (Please show how you worked out the problem)

OpenStudy (dogzcatz):

So it is 3a/1 = 10/3. Now multiply by 3. 9a = 10 a = 10/9

OpenStudy (anonymous):

Thank you so much!